# Berkovich Motives

### v1

@@ 9. $\mathcal{D}_{\mathrm{mot}}(X)$
For any map $f:Y=\mathcal{M}_{\mathrm{arc}}(B)\to X=\mathcal{M}_{\mathrm{arc}}(A)$ of finite cohomological dimension of affine Berkovich spaces, the functor
@@
We already had colimit-preserving right adjoints $f_{\ast}$ on $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}$ satisfying base change. They pass on to $\mathcal{D}_{\mathrm{mot}}$, taking a collection
@@
to the collection $f_{\ast}N=(M_{n})_{n}$ with $M_{n}=f_{\ast}N_{n}$; this obviously still commutes with colimits and satisfies base change. We have to see that it satisfies the projection formula. Checking isomorphisms at stalks, we can assume that $X=\mathcal{M}_{\mathrm{arc}}(C)$ with $C$ a non-discrete algebraically closed Banach field. The case $C=\mathbb{C}$ is trivial, so we can assume that $C$ is nonarchimedean.
We can write $Y$ as a limit of closed subsets of finite-dimensional balls over $C$, which writes $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(Y)$ and $\mathcal{D}_{\mathrm{mot}}(Y)$ as a corresponding colimit of presentable stable $\infty$-categories. We can thus reduce to the case that $Y$ is a closed subset of some finite-dimensional ball. One can then further reduce to the case that $Y$ is a finite-dimensional ball (as the result is clear for closed immersions). By induction, we can reduce to the case of a $1$-dimensional ball. This can be embedded into $\mathbb{P}^{1}_{C}$. Now the result follows from Proposition 9.3 below. ∎
@@
It suffices to check the result for objects in the image of $j_{!}$ where $j:\mathbb{A}^{1}_{X}\to\mathbb{P}^{1}_{X}$ is the open immersion (it is then also the case fo other points at infinity, and these generate). We write $f^{\circ}_{!}$ for $f_{\ast}j_{!}$; after we completed this proof, this is the !!!-functor for $f^{\circ}:\mathbb{A}^{1}_{X}\to X$. It suffices to see that the resulting composite
@@
In particular, starting with the category of affine arc-sheaves, and the class of maps of finite cohomological dimension as the !!!-able (and proper) maps, we get a $6$-functor formalism $X\mapsto\mathcal{D}_{\mathrm{mot}}(X)$. By general extension procedures, cf. [HM24], we get a $6$-functor formalism $X\mapsto\mathcal{D}_{\mathrm{mot}}(X)$ on all small arc-stacks, for a certain class of !!!-able maps of small arc-stacks.
@@
Open immersions $j:U\hookrightarrow X$ are !!!-able and $j_{!}$ is the left adjoint of pullback $j^{\ast}$.
@@
We can check this for $X=\mathcal{M}_{\mathrm{arc}}(A)$ affine. It suffices to cover $U$ by !!!-able open subsets with the desired property; then it follows for $U$ by gluing. So we can assume $U=\{|f|<1\}$ for some $f\in A$. This is the union of the affine subsets $\{|f|\leq c\}$ over $c<1$, and these form a !!!-cover of $U$. It follows that $U$ is !!!-able, and it is a standard exercise to compute the resulting $j_{!}$. (Indeed, it is colimit-preserving and has the correct restriction to all $\{|f|\leq c\}$.) ∎
@@
All étale maps (i.e. finitary maps with finite geometric fibres) are !!!-able and cohomologically étale.
@@
Any smooth space is locally étale over an affine space, reducing us to the previous result. ∎
@@
Indeed, the arc-sites of $X$ and $X^{\flat}$ have equivalent bases given by affinoid perfectoids, hence their arc-topoi are equivalent, and the resulting equivalence passes through all further steps of the construction. ∎
@@ 10. $\mathcal{D}_{\mathrm{mot}}(C)$
By Proposition 9.7, we can assume that $C$ has equal characteristic. Pick a splitting $k\to C$ of its residue field. It is clear that $\mathbb{Z}_{\mathrm{mot}}[Y](-j)$ is dualizable when $Y$ is smooth projective over $C$, in particular this happens for $Y=X_{C}$. Also, the unit is compact, so all dualizable objects are compact. For the rest, it suffices to see that the $\mathbb{Z}_{\mathrm{mot}}[X_{C}](-j)$ generate $\mathcal{D}_{\mathrm{mot}}(C)$. Already $\mathbb{Z}$ itself generates the torsion part, so we can work with rational coefficients. Assume that $M\in\mathcal{D}_{\mathrm{mot}}(C)$ has the property that it has no maps from $\mathbb{Z}_{\mathrm{mot}}[X_{C}](-j)$ for all such $X$ and $j$; we need to show that $M=0$. It suffices to show that $M(C^{\prime})=0$ for all complete algebraically closed extensions $C^{\prime}/C$. We argue by induction on the topological transcendence degree. If $C_{d}/C$ is some extension of topological transcendence degree $d$, pick some $C_{d-1}\subset C_{d}$ of topological transcendence degree $d-1$ such that $C_{d}/C_{d-1}$ is of transcendence degree $1$. It then corresponds to one of the types of points (2), (3) or (4) on the Berkovich line over $C_{d-1}$. But type (4) points are intersections of balls, so evaluation at $C_{d}$ agrees with evaluation at $C_{d-1}$, giving the result by induction. Type (3) points are intersections of annuli, so evaluation at $C_{d}$ agrees with evaluation at $C_{d-1}$ of Tate twists, so again we get the result by induction. Thus, only type (2) remains. But if the residue $k_{d}$ of $C_{d}$ has transcendence degree $<d$ over the residue field $k$ of $C$, we could always choose $C_{d-1}$ to already have residue field $k_{d-1}=k_{d}$, thereby avoiding type (2). Thus, only the case that $k_{d}$ has transcendence degree $d$ is relevant. In that case, we can write $k_{d}$ as a filtered colimit of affine smooth $k$-algebras $R_{i}$ such that $\mathrm{Spec}(R_{i})\subset X_{i}$ has a smooth projective compactification $X_{i}$ whose boundary is a normal crossing divisor. Now the motive of $R_{i}\hat{\otimes}C$ can be described in terms of the motive of $X_{i,C}$, and lower-dimensional varieties, by induction; and the completed colimit of the $R_{i}\hat{\otimes}C$ is $C_{d}$. ∎
@@
Let $f:Y\to X$ be a proper 00-truncated map of locally finite cohomological dimension between small arc-stacks. Then
@@
We can then rerun the construction of $6$-functor formalisms using this class of proper maps, to get a more general class of !!!-able maps.
@@
The unit is compact by assumption on finite cohomological dimension. First we check dualizability; equivalently, being compactly assembled. For this, it suffices to exhibit sufficiently many compact maps. But whenever $U\subset V\subset\mathbb{A}^{n}_{X}$ is an inclusion of open subsets that factors over a compact subset $K\subset\mathbb{A}^{n}_{X}$, then the map $\mathbb{Z}_{\mathrm{mot}}[U]\to\mathbb{Z}_{\mathrm{mot}}[V]$ is compact, as the induced map $\mathrm{Hom}(V,-)\to\mathrm{Hom}(U,-)$ factors over evaluation at $K$, and evaluation at $K$ commutes with filtered colimits. As any $V\subset\mathbb{A}^{n}_{X}$ can be written as a union of such $U$ strictly contained in $V$, we see that $\mathbb{Z}_{\mathrm{mot}}[V]$ can be written as a filtered colimit with compact transition maps. To deduce dualizability of $\mathcal{D}_{\mathrm{mot}}(X)$, it remains to see that those objects generate $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ (then their negative Tate twists generate all of $\mathcal{D}_{\mathrm{mot}}(X)$).
Let $\mathcal{C}\subset\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ be the subcategory generated $\mathbb{Z}_{\mathrm{mot}}[U]$ where $U\subset\mathbb{A}^{n}_{X}$ is open. By the preceding paragraph, we see that $\mathcal{C}$ is compactly assembled and that the inclusion preserves compact morphisms, and hence its right adjoint preserves colimits. In particular, its formation commutes with passing to stalks of $|X|$, and so the claim $\mathcal{C}=\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ can be reduced to the case that $|X|$ is a point. Then $X=\mathcal{M}_{\mathrm{arc}}(C)/G$ for some profinite group $G$ acting continuously on an algebraically closed non-discrete Banach field $C$. Now the result can be proved in the same way as Proposition 5.13.
@@
In general, it suffices to see that there is a generating family of objects that are sequential colimits along trace-class maps. Pick any $x\in X$ and write it as a cofiltered limit of closed neighborhoods $Z_{i}\subset X$. Using Lemma 10.5 again, the compact dualizable objects in the fibre at $x$ spread to some $Z_{i}$. If $M\in\mathcal{D}_{\mathrm{mot}}(Z_{i})$ is compact dualizable and $U_{i}\subset Z_{i}$ is an open neighborhood of $x$ in $X$, without loss of generality a countable union of closed subsets, then the extension by 00 from $U_{i}$ to $X$ of $M|_{U_{i}}$ is such a sequential colimit along trace-class maps, and these objects generate. ∎
@@
The map $f$ has cohomological dimension 00, so $f_{!}=f_{\ast}$ commutes with colimits and satisfies the projection formula. Moreover, $f_{\ast}$ is conservative: This can be checked after passing to geometric fibres, but when $X$ is a geometric point, then $Y$ is a profinite set times a geometric point, where global sections are conservative. Now the first result follows from Barr–Beck–Lurie. The colimit statement is a consequence of $f_{\ast}\mathbb{Z}=\mathrm{colim}_{i}f_{i\ast}\mathbb{Z}$ and the behaviour of module categories under colimits. ∎
@@
Let $X\to S\leftarrow Y$ be a diagram of qcqs arc-stacks as in Proposition 10.3. Then the exterior tensor product defines a fully faithful functor
@@
Relative tensor products of rigid categories remain rigid, and for any symmetric monoidal functor of rigid categories, the right adjoint functor is colimit-preserving and linear over the base category. Fully faithfulness is then equivalent to showing that the right adjoint applied to the unit yields the unit. But this follows from the Künneth formula which follows formally from the projection formula. ∎

### v2

@@ 9. $\mathcal{D}_{\mathrm{mot}}(X)$
For any map $f:Y=\mathcal{M}_{\mathrm{arc}}(B)\to X=\mathcal{M}_{\mathrm{arc}}(A)$ of finite cohomological dimension of affine Berkovich spaces such that $Y$ admits an embedding into a finite-dimensional ball over $X$, the functor
@@
Using descent (and [Lur17, Corollary 4.7.4.18]), one can formally deduce that the similar result holds true more generally when $f$ is only arc-locally (on the target) of this form. An even more general form of the result will be proved in Corollary 10.2 below.
@@
We already had colimit-preserving right adjoints $f_{\ast}$ on $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}$ satisfying base change by Theorem 4.20. They pass on to $\mathcal{D}_{\mathrm{mot}}$, taking a collection
@@
to the collection $f_{\ast}N=(M_{n})_{n}$ with $M_{n}=f_{\ast}N_{n}$; this obviously still commutes with colimits and satisfies base change. We have to see that it satisfies the projection formula. Checking isomorphisms at stalks, we can assume that $X=\mathcal{M}_{\mathrm{arc}}(C)$ with $C$ a non-discrete algebraically closed Banach field. The case $C=\mathbb{C}$ is trivial, using Theorem 4.14 and the usual projection formula in topology, so we can assume that $C$ is nonarchimedean.
One can then further reduce to the case that $Y$ is a finite-dimensional ball (as the result is clear for closed immersions, e.g. by the final part of Theorem 4.20). By induction, we can reduce to the case of a $1$-dimensional ball. This can be embedded into $\mathbb{P}^{1}_{C}$. Now the result follows from Proposition 9.3 below. ∎
@@
It suffices to check the result for objects in the image of $j_{!}$ where $j:\mathbb{A}^{1}_{X}\to\mathbb{P}^{1}_{X}$ is the open immersion (it is then also the case fo other points at infinity, and these generate). We write $f^{\circ}_{!}$ for $f_{\ast}j_{!}$; after we completed this proof, this is the !!-functor for $f^{\circ}:\mathbb{A}^{1}_{X}\to X$. It suffices to see that the resulting composite
@@
For the following discussion of six-functor formalisms, we make extensive use of the references [Sch25b] and [HM24]. Starting with the category of affine arc-sheaves, and the class of maps of finite cohomological dimension as the !!-able (and proper) maps, we get a $6$-functor formalism $X\mapsto\mathcal{D}_{\mathrm{mot}}(X)$. By general extension procedures, cf. [Sch25b, Theorem 5.19] and [HM24, Theorem 3.4.11], we get a $6$-functor formalism $X\mapsto\mathcal{D}_{\mathrm{mot}}(X)$ on all small arc-stacks, for a certain class of !!-able maps of small arc-stacks.
@@
Open immersions $j:U\hookrightarrow X$ are !!-able and $j_{!}$ is the left adjoint of pullback $j^{\ast}$.
@@
We can check this for $X=\mathcal{M}_{\mathrm{arc}}(A)$ affine. It suffices to cover $U$ by !!-able open subsets with the desired property; then it follows for $U$ by gluing. So we can assume $U=\{|f|<1\}$ for some $f\in A$. This is the union of the affine subsets $\{|f|\leq c\}$ over $c<1$, and these form a !!-cover of $U$. It follows that $U$ is !!-able, and it is a standard exercise to compute the resulting $j_{!}$. (Indeed, it is colimit-preserving and has the correct restriction to all $\{|f|\leq c\}$.) ∎
@@
All étale maps (i.e. finitary maps with finite geometric fibres) are !!-able and cohomologically étale.
@@
In the following corollary, when working over a Banach field $K$, we use Berkovich’s notion of smooth $K$-analytic spaces. These correspond to smooth separated rigid-analytic spaces over $K$.
@@
Any smooth space is locally étale over $\mathbb{A}^{n}$, reducing us to the previous result (noting that boundaryless étale maps of $K$-Berkovich spaces are étale in our sense; this reduces to the case of finite étale maps and open immersions, which are clear). ∎
@@
We already proved the effective version in Proposition 6.8. Inverting the Tate twist (which is preserved under tilting), the result follows. ∎
@@ 10. $\mathcal{D}_{\mathrm{mot}}(C)$
By Proposition 9.7, we can assume that $C$ has equal characteristic. Pick a splitting $k\to C$ of its residue field. As in any six-functor formalism, pushforward along cohomologically smooth and proper maps preserves dualizable objects, it is clear that $\mathbb{Z}_{\mathrm{mot}}[Y](-j)$ is dualizable when $Y$ is smooth projective over $C$, in particular this happens for $Y=X_{C}$. Also, the unit is compact, so all dualizable objects are compact. For the rest, it suffices to see that the $\mathbb{Z}_{\mathrm{mot}}[X_{C}](-j)$ generate $\mathcal{D}_{\mathrm{mot}}(C)$. Already $\mathbb{Z}$ itself generates the torsion part, so we can work with rational coefficients. Assume that $M\in\mathcal{D}_{\mathrm{mot}}(C)$ has the property that it has no nonzero maps from $\mathbb{Z}_{\mathrm{mot}}[X_{C}](-j)$ for all such $X$ and $j$; we need to show that $M=0$. It suffices to show that $M(C^{\prime})=0$ for all complete algebraically closed extensions $C^{\prime}/C$. We argue by induction on the topological transcendence degree. If $C_{d}/C$ is some extension of topological transcendence degree $d$, pick some $C_{d-1}\subset C_{d}$ of topological transcendence degree d1d-1 such that $C_{d}/C_{d-1}$ is of transcendence degree $1$. It then corresponds to a completed algebraic closure of one of the types of points (2), (3) or (4) on the Berkovich line over $C_{d-1}$. The residue field at the point may not be algebraically closed, but any of its finite extensions can also be realized by a point of the same type in the Berkovich line over $C_{d-1}$. Type (4) points are intersections of balls, so evaluation at $C_{d}$ agrees with evaluation at $C_{d-1}$, giving the result by induction. Type (3) points are intersections of annuli, so evaluation at $C_{d}$ agrees with evaluation at $C_{d-1}$ of Tate twists, so again we get the result by induction. Thus, only type (2) remains. But if the residue $k_{d}$ of $C_{d}$ has transcendence degree <d<d over the residue field $k$ of $C$, we could always choose $C_{d-1}$ to already have residue field $k_{d-1}=k_{d}$, thereby avoiding type (2). Thus, only the case that $k_{d}$ has transcendence degree $d$ is relevant. Write $\mathcal{O}_{C_{d}}$ as a completed filtered colimit of algebras topologically of finite type $R_{i}$ over $\mathcal{O}_{C}$. These algebras are generically smooth, so we can assume that the $R_{i}$ are smooth. We can then find an isomorphism $R_{i}\cong\overline{R}_{i}\hat{\otimes}_{k}\mathcal{O}_{C}$ for smooth $k$-algebras $\overline{R}_{i}$. Using alterations, we can arrange that $\mathrm{Spec}(\overline{R}_{i})\subset X_{i}$ has a smooth projective compactification $X_{i}$ whose boundary is a normal crossing divisor. Now the motive of $\overline{R}_{i}\hat{\otimes}_{k}C$ can be described in terms of the motive of $X_{i,C}$, and lower-dimensional varieties, by induction; and the completed colimit of the $\overline{R}_{i}\hat{\otimes}_{k}C$ is $C_{d}$ by construction. ∎
@@
Let $f:Y\to X$ be a proper 0-truncated map of locally finite cohomological dimension between small arc-stacks. Then
@@
We can then rerun the construction of $6$-functor formalisms using this class of proper maps (still between affine arc-stacks) [^5] In [HM24, Theorem 3.4.11], the resulting class of !!-able maps depends on the choice of basic geometric objects, in this case affine arc-stacks: Namely, to check that a map of stacks is !!-able, it suffices to check this after pullback to each affine arc-stack. If we removed the affineness condition here, the resulting condition would be harder to check in practice., to get a more general class of !!-able maps.
@@
The underlying topological space $|X|$ is defined via left Kan extension from the affine case. If $X$ is qcqs, then choosing a surjection $\mathcal{M}_{\mathrm{arc}}(A)\to X$ and $\mathcal{M}_{\mathrm{arc}}(B)$ surjecting onto the equivalence relation, we can write $|X|$ as the quotient $\mathcal{M}(A)/\mathcal{M}(B)$. This is a quotient of a compact Hausdorff space by a closed equivalence relation and hence itself compact Hausdorff. Similarly, the underlying topological space of $\mathbb{A}^{n}_{X}$ is a locally compact Hausdorff space.
@@
The unit is compact by assumption on finite cohomological dimension. First we check dualizability; equivalently, being compactly assembled. For this, it suffices to exhibit sufficiently many compact maps. But whenever $U\subset V\subset\mathbb{A}^{n}_{X}$ is an inclusion of open subsets that factors over a quasicompact subset $K\subset\mathbb{A}^{n}_{X}$, then the map $\mathbb{Z}_{\mathrm{mot}}[U]\to\mathbb{Z}_{\mathrm{mot}}[V]$ is compact, as the induced map $\mathrm{Hom}(V,-)\to\mathrm{Hom}(U,-)$ factors over evaluation at $K$, and evaluation at $K$ commutes with filtered colimits. As any $V\subset\mathbb{A}^{n}_{X}$ can be written as a union of such $U$ strictly contained in $V$, we see that $\mathbb{Z}_{\mathrm{mot}}[V]$ can be written as a filtered colimit with compact transition maps. To deduce dualizability of $\mathcal{D}_{\mathrm{mot}}(X)$, it remains to see that those objects generate $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ (then their negative Tate twists generate all of $\mathcal{D}_{\mathrm{mot}}(X)$).
Let $\mathcal{C}\subset\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ be the subcategory generated $\mathbb{Z}_{\mathrm{mot}}[U]$ where $U\subset\mathbb{A}^{n}_{X}$ is open. By the preceding paragraph, we see that $\mathcal{C}$ is compactly assembled and that the inclusion preserves compact morphisms, and hence its right adjoint preserves colimits. In particular, its formation commutes with passing to stalks of $|X|$, and so the claim $\mathcal{C}=\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)$ can be reduced to the case that $|X|$ is a point. Then $X=\mathcal{M}_{\mathrm{arc}}(C)/G$ for some profinite group $G$ acting continuously on an algebraically closed non-discrete Banach field $C$ (see [Sch17, Proposition 21.9] for the argument). Now the result can be proved in the same way as Proposition 5.13, using $G$-equivariant open subsets of $\mathbb{A}^{n}_{C}$. The points of $\mathbb{A}^{n}_{C}/G$ are then of the form $\mathcal{M}_{\mathrm{arc}}(C^{\prime})/H$ for a profinite group $H$ acting faithfully on complete algebraically closed nonarchimedean fields $C^{\prime}$, and the arguments still apply in this case, replacing finite extension fields of Banach fields by open subgroups of $H$.
@@
In general, it suffices to see that there is a generating family of objects that are sequential colimits along trace-class maps, cf. [Ram24b, Corollary 4.57]. Pick any $x\in X$ and write it as a cofiltered limit of closed neighborhoods $Z_{i}\subset X$. Using Lemma 10.5 again, the compact dualizable objects in the fibre at $x$ spread to some $Z_{i}$. If $M\in\mathcal{D}_{\mathrm{mot}}(Z_{i})$ is compact dualizable and $U_{i}\subset Z_{i}$ is an open neighborhood of $x$ in $X$, without loss of generality a countable union of closed subsets, then the extension by 0 from $U_{i}$ to $X$ of $M|_{U_{i}}$ is such a sequential colimit along trace-class maps, and these objects generate. ∎
@@
The map $f$ has cohomological dimension 0, so $f_{!}=f_{\ast}$ commutes with colimits and satisfies the projection formula. Moreover, $f_{\ast}$ is conservative: This can be checked after passing to geometric fibres, but when $X$ is a geometric point, then $Y$ is a profinite set times a geometric point, where global sections are conservative. Now the first result follows from Barr–Beck–Lurie [Lur17, Theorem 4.7.3.5]. The colimit statement is a consequence of $f_{\ast}\mathbb{Z}=\mathrm{colim}_{i}f_{i\ast}\mathbb{Z}$ and the behaviour of module categories under colimits. ∎
@@
Let $X\xrightarrow{f}S\xleftarrow{g}Y$ be a diagram of qcqs arc-stacks as in Proposition 10.3. Then the exterior tensor product defines a fully faithful functor
@@
Relative tensor products of rigid categories remain rigid, and for any symmetric monoidal functor of rigid categories, the right adjoint functor is colimit-preserving and linear over the base category. It follows that it suffices to check that the unit of the adjunction becomes an equivalence after evaluation at the unit, or the slightly more refined statement that it holds after pushforward to $\mathcal{D}_{\mathrm{mot}}(S)$. This means that we have to check that for all $M\in\mathcal{D}_{\mathrm{mot}}(X)$ and $N\in\mathcal{D}_{\mathrm{mot}}(Y)$ with associated $M\boxtimes N\in\mathcal{D}_{\mathrm{mot}}(X\times_{S}Y)$, the natural map
$f_{\ast}M\otimes g_{\ast}N\to(f\times_{S}g)_{\ast}(M\boxtimes N)$
is an isomorphism. But this is the usual Künneth formula which holds in any six-functor formalism (as a consequence of the projection formula). ∎